OpenStudy (cookiemonster18):

Evaluate X/Y for X=4/3 and Y=6/11.

OpenStudy (anonymous):

22/9

jabez177 (jabez177):

Basically, \[\frac{ 4 }{ 3 }\div \frac{ 6 }{ 11 }\]

OpenStudy (cookiemonster18):

how @CharlieJay

jabez177 (jabez177):

So whenever we are dividing fractions, we can also multiply the reciprocate.

jabez177 (jabez177):

\[\frac{ 4 }{ 3 }\times \frac{ 11 }{ 6 }\]

OpenStudy (anonymous):

Cross multiplying is the way I do it in short. 4*11/3*6 = 44/18 = 22/9

jabez177 (jabez177):

Now what is 4 x 11?

OpenStudy (cookiemonster18):

I did that in my calculator and it gave it to me in decimals then I did it in a different one and it said error :/.... but thank you @CharlieJay and @jabez177

jabez177 (jabez177):

@CharlieJay, welcome to OpenStudy! We are supposed to guide the person to the answer, not just give it out. :)

OpenStudy (cookiemonster18):

4x11 is 44

jabez177 (jabez177):

Good, now what is 3 x 6?

OpenStudy (cookiemonster18):

3x6 is 18

jabez177 (jabez177):

\[\frac{ 44 }{ 18 }\div \frac{ 2 }{ 2 }\]

jabez177 (jabez177):

What is 44 divided by 2?

OpenStudy (cookiemonster18):

22

jabez177 (jabez177):

Now what is 18 divided by 2?

OpenStudy (cookiemonster18):

9

jabez177 (jabez177):

Perfect, when you put that in a fraction it comes out as\[\frac{ 22 }{ 9}\]

jabez177 (jabez177):

And that is how you get your answer. :) It is in simplest form. :)

OpenStudy (anonymous):

@jabez177 Sorry about that, I guess I'll type out the longer way just for reference. In order to divide 2 fractions similar to these you would need to cross multiply taking the first fractions numerator and multiply it by the second fractions denominator and that will give you your new numerator and then you take the second fractions numerator and multiply it by the first fractions denominator and that will give you your new denominator and then from there you simplify. You are also just multiplying by the reciprocal which is as simple as flipping the fractions around making the numerator the denominator and the denominator the numerator. 4/3 / 6/11 4/3 * 11/6 4*11 = 44 3*6 = 18 New fraction is equal to 44/18 and that simplified is 22/9. Once again sorry about that.

jabez177 (jabez177):

It's okay @CharlieJay, welcome to OpenStudy! :)